OpenStudy (anonymous):

I don't know where to start with this question. An air tanker is helping fight a forest fire in a mountainous region. The air tanker fills by skimming a lake, flies with the wind to the fire, empties, and flies against the wind to return to the lake to refill. If the plane flies at 294 miles per hour in still air, the wind speed is 14 miles per hour, and one round trip from the lake to the fire and back to the lake takes at most 90 minutes, how far, at most, is the fire from the lake?

OpenStudy (perl):

lets use the equation distance = rate * time

OpenStudy (perl):

flying from the lake to the fire : distance = (294 + 14) * t1 flying from the fire to the lake: distance = (294 - 14) * t2

OpenStudy (perl):

d / (294 + 14) = t1 d/ (294 - 14) = t2 we are given that t1 + t2 <= 90/60 hours

OpenStudy (perl):

d/308 + d/280 <= 90/60 solve this inequality for d , in miles